If the surface area of a cube is 24 m2, find the length of one side
Found 2 solutions by edjones, rapaljer:
Answer by edjones(7569)

24/6=4 m^2 area of one side
sqrt(4)=2m length of one side.
.
Ed

Answer by rapaljer(4667)

There are 6 faces of a cube. Each one is a square. If the side is x, then the total surface area is 6x^2 = 24.

6x^2=24
x^2=4
x=2 meters

Dr. Robert J. Rapalje, Retired